OpenStudy (anonymous):

A positron and electron are released from a distance L away from each other, at what time t will they meet? Use conservation laws. Any help, greatly appreciated.

OpenStudy (anonymous):

What have you go so far?

OpenStudy (anonymous):

\[\frac{kq^2}{x}=mv^2+\frac{kq^2}{L}\]

OpenStudy (anonymous):

You know how people always tell you to elaborate what your symbols are supposed to mean. And how you got your equations - whats the idea behind them and so on... listen to them ;)

OpenStudy (anonymous):

total energy = kinetic + potential. when I work this stuff out I end up with \[dt = \frac{\sqrt {m}dr}{2a \sqrt{k} \sqrt{\frac{1}{L}-\frac{1}{r}}} \] I get stuck at this point because of the integral. However, I was wondering if anyone knew if the initial equation I used is correct, and if it is, then my second question is: how do you go about integrating.

OpenStudy (anonymous):

(the a is suppose to be an L)

OpenStudy (anonymous):

Okay ahm.. why _will_ they meet in the first place? They are released a distance apart from each other, so why would they meet?

OpenStudy (anonymous):

there's an attractive force, but the professor asked us to use conservation laws

OpenStudy (anonymous):

1st of all .. as I told you above, you need to say what you mean.. WHY is there and attractive force? Has it got a name? How does it work? etc, etc.. 2nd: what's the "but" doing there? Why do you think a force is something evil when talking about conservation of... whatever.. ??

OpenStudy (anonymous):

raccoondog, in your formula relating dt to dr , if you let r=L you get a zero in the denominator. This is not good. I suspect you made a mistake somewhere.